OpenStudy (anonymous):

Will give medal and fan.. (: What exponential function is the best fit for the data in the table? x f(x) 1 -4 3 -1 4 3

OpenStudy (anonymous):

f(x) = 4(3)x - 1 + 4 f(x) = 4(3)x - 1 - 4 f(x) = one fourth(3)x - 1 + 4 f(x) = one fourth(3)x - 1 - 4

OpenStudy (anonymous):

@hartnn .... sorry I keep asking you for help today xp You just explain things in a way that I understand it... other people do not.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

I think it might be B